### 19 Dec 2023 v1

# ProS: Prompting-to-simulate Generalized knowledge for Universal Cross-Domain Retrieval

Abstract
The goal of Universal Cross-Domain Retrieval (UCDR) is to achieve robust performance in generalized test scenarios, wherein data may belong to strictly unknown domains and categories during training. Recently, pre-trained models with prompt tuning have shown strong generalization capabilities and attained noteworthy achievements in various downstream tasks, such as few-shot learning and video-text retrieval. However, applying them directly to UCDR may not sufficiently to handle both domain shift (i.e., adapting to unfamiliar domains) and semantic shift (i.e., transferring to unknown categories). To this end, we propose Prompting-to-Simulate (ProS), the first method to apply prompt tuning for UCDR. ProS employs a two-step process to simulate Content-aware Dynamic Prompts (CaDP) which can impact models to produce generalized features for UCDR. Concretely, in Prompt Units Learning stage, we introduce two Prompt Units to individually capture domain and semantic knowledge in a mask-and-align way. Then, in Context-aware Simulator Learning stage, we train a Content-aware Prompt Simulator under a simulated test scenarios to produce the corresponding CaDP. Extensive experiments conducted on three benchmark datasets show that our method achieves new state-of-the-art performance without bringing excessive parameters. Our method is publicly available at https://anonymous.4open.science/r/ProS
1 Introduction
[…]
Cross-Domain Retrieval (CDR) [43, 11] is an important task in Information Retrieval (IR) systems that utilizes data from one domain (e.g. Infograph, Sketch) as a query to search semantically similar examples in another domain (e.g. natural image). Despite achieving promising results, existing methods [34, 11, 2, 4] heavily follow a close-set learning setting, where domains and categories in both training and testing data are pre-defined. Such evaluation limiting the model to the training samples and unable to tackle the open-set applications such as e-commerce search [6, 10, 25] and recommendation [18, 10], where searched product not only comes from various new domains (e.g. styles) but is also accompanied by the emergence of new categories. Towards this goal, one emerging research direction is proposed, namely Universal Cross-Domain Retrieval (UCDR) [29], shown in Fig. 1(a), for performing well under generalized test scenarios, where test data may belong to strictly unknown domains and categories.
The core challenge for UCDR is how to empower models with the capabilities of both domain shift (i.e. adapting to unfamiliar domains) and semantic shift (i.e. transferring to unknown categories). The mainstream approaches [29, 37] focus on fine-tuning visual models, e.g. ResNet [14] and ViT [8], with the guidance of expert knowledge from semantic information or pre-trained models. Compared with ResNet, ViT has stronger representational capability and contains more knowledge gained from larger pre-training data, which facilitates the generalization of the model for unknown images. This raises a intuitive question: does the extensive general knowledge inherent in large-scale pre-trained models can enhance it’s ability to tackle UCDR task where priori information about test samples are strict unknown?
With this question in mind, we empirically conduct a confirmatory experiment on the widely-recognized large-scale pre-trained model, CLIP [31] under UCDR. Specifically, we full fine-tune the CLIP and compare its performance with ViT and ResNet, as shown in Fig. 1 (b). From the figure, we can see that CLIP yields tremendous gains, reaching a 22.27% improvement in mAP@200 compared with ViT. This result proves the effectiveness of common knowledge inherent in large-scale pre-trained CLIP for UCDR task.
However, the full fine-tuning strategy inevitably forgets the useful knowledge gained in the large-scale pre-training phase [19]. Besides, since the entire model is updated, this strategy brings significant computational costs. To alleviate above problems, we take the first step to leverage prompt tuning [41, 17] which is a more flexible and lightweight strategy than full fine-tuning. However, when directly applying existing prompt tuning methods [1, 17, 41, 42] to UCDR, i.e. $\triangle$ in Fig. 1 (b), we find that these methods show a fine benefit in terms of computational cost without a strong advantage or even worse over full fine-tuning CLIP in terms of mAP@200. This may be attributed that these methods do not fulfill the properties of UCDR, contributing to poor generalization. Therefore, we pose a question: how to design a more effective prompt tuning method for the UCDR task?
Towards above question, we propose a novel prompt tuning method in a simulated way, named Prompting-to-Simulate (ProS), to effectively leverage prompt tuning to mine the generalized knowledge from CLIP for UCDR. Specifically, ProS introduces a two-stage learning to simulate two content-aware dynamic prompts (CaDP) which reflect input sample’s domain and category, respectively. In Prompt Unit Learning (PUL), two groups of distinct learnable prompts are established: domain prompt units and semantic prompt units. These two units are employed to extract domain knowledge and semantic knowledge from source data in a mask-and-align strategy. Then, in Context-aware Simulator learning (CSL) , we train a Content-aware Prompt Simulator (CaPS) under simulated test scenarios to dynamically generate content-aware dynamic prompts (CaDP) based on the learned prompt units. With the CaDP, we can impact CLIP’s knowledge to gain more generalizable representation and achieve better performance compared with existing prompt tuning methods in UCDR, as shown in Fig. 1 (b). In addition, our method uses considerably fewer learnable parameters compared to full fine-tuning, accounting for only 12.84% of the model parameters, and does not introduce excessive learnable parameters compared to prompt tuning methods.
Our main contributions can be summed as follows:
  • We are the first to investigate how to adapt CLIP with prompt for UCDR.
  • We propose a prompt-based method named Prompting-to-Simulate (ProS), which learns the generalized knowledge to deal with open-set scenarios.
  • Extensive experiments on three benchmark datasets show that our ProS achieves new state-of-the-art results compared with prompt-based methods without bringing excessive parameters.
2 Related Work
2.1 Universal Cross-Domain Retrieval
UCDR requires solving domain and category generalization simultaneously which can be roughly regarded as a combination of Domain Generalization (DG) [24, 38] and Zero-Shot Learning (ZSL) [39, 13]. In DG, some approaches synthesize unseen samples using data augmentation strategy [28] or generative adversarial networks (GANs) [40] to handle the domain shift. In ZSL, some researchers leverage the auxiliary data, including human-annotated attribute information  [21, 20], text description[32] or knowledge graph [22] to learn relationship between seen categories and pre-define unseen categories names. It should be clarified that UCDR does not require pre-define unseen class names. Inspired by above approaches, SnMpNet [29] trains the model with the mix-up data augmentation strategy and align features with semantic information. Additionally, SASA[37] deploys a ViT-based model and train categorical prototypes to handle the semantic shift. Different from these works, we make the first attempt to apply a large-scale pre-trained model, i.e. CLIP, to the UCDR task, and proposed a new prompt tuning method to improve the generalization ability of the CLIP.
2.2 Vision-Language Pre-training Models
Motivated by the success in self-supervised learning, Vision-Language Pre-training has emerged as a prominent topic. CLIP [31] initially demonstrated that equipped with large-scale image-text pairs, contrastive learning framework can achieve a performance comparable to fully supervised baselines. Subsequently, ALIGN [16] scale the training dataset to billions and gain a better vision-language representation. Another line of works [12, 5] leverage pre-trained object detection models like Faster-RCNN [33] to extract image regional features offline for training multi-modal transformers. These models show impressive performance on various vision-language tasks, as they are capable of processing both single-modal and multi-modal inputs. In this paper, we make the first attempt to leverage the powerful CLIP for the UCDR task.
2.3 Prompt Tuning
Prompt Tuning [27], as a new paradigm, initially surface in NLP for adapting pre-trained language models (PLMs) [7, 3] to downstream tasks. These prompts can either be handcrafted for specific tasks [36] or learned automatically via gradients which termed as “Prompt Learning” [23]. Driven by the success of prompts paradigm in NLP, various studies like CoOP [41] and CoCoOp [42] apply text prompting in multi-modal scenarios, while VPT [17] and VP [1] introduces these paradigm into vision transformer. Despite the pioneering successes of VP and VPT in visual models, we find that they are static after training which is not flexible to handle the generalized scenarios. In contrast, our method propose a Prompt-to-Simulate method to well address UCDR task.
3 Method
Our method makes the first attempt to apply CLIP for Universal Cross-Domain Retrieval (UCDR) with a novel prompt learning method named ProS, as shown in Fig. 2. We begin with a brief overview of some important preliminaries about UCDR, CLIP and prompt tuning paradigm (Sec. 3.1). Then, a detailed description of our proposed Prompting-to-Simulate (Pros) is provided (Sec. 3.2), including two training stages: Prompt Unit Learning (PUL) and Context-aware Simulator Learning (CSL). Finally, we will introduce how to perform retrieval using our proposed method (Sec. 3.3).
Figure 2: Overview of our proposed ProS. In Prompt Units Learning Stage, we capture knowledge from source data into domain prompts units $DP$ and semantic prompts units $SP$ by masking irrelevance prompts. In Context-aware Prompt Simulation Stage, we train a Context-aware Prompt Simulator (CaPS) with a mask operation to dynamically convey prompt templates $PT$ to two Content-aware Dynamic Prompts (CaDP) to simulate unknown domains and categories. In retrieval phase, we employ CaPS to produce CaDP which impact the CLIP image encoder to convert unseen samples into suitable embeddings for retrieval. The gray parts indicate masked prompts.
3.1 Preliminary
Problem Formulation. Universal Cross-Domain Retrieval (UCDR) [29] aims to accurately retrieve images with queries from unseen domains and classes. From a general perspective, we assume that training dataset is composed of $K$-different ($K\geq 2$) source domains (Real, Sketch, Quickdraw, etc.) which can be formulated as $\mathcal{D}_{train}=\bigcup_{d\in\left\{1,\ldots,K\right\}}\left\{\boldsymbol{x}_{i}^{d},y^{tr}_{i}\right\}_{i=1}^{N_{d}}$. The $d^{th}$ domain has $N_{d}$ images, where each image $\boldsymbol{x}_{i}^{d}$ is belong to class $y^{tr}_{i}\in\mathcal{C}_{train}$. In the testing phase, the gallery set and query set are needed to enable image retrieval. Specifically, query set $\mathcal{D}_{query}=\left\{\boldsymbol{x}_{i}^{q},y^{te}_{i}\right\}_{i=1}^{N_{q}}$ consists of $N_{q}$ samples belong to unseen class $y_{i}^{\text{te}}\in\mathcal{C}_{\text{test}}$ from the query domain $q$, while gallery set consists of real images. To simulate a more realistic retrieval scenario, the gallery set has two settings: 1) all samples belong to unseen class, termed Unseen Gallery; (2) samples belong to both seen class and unseen class, termed Mixed Gallery.
The UCDR protocol can be separated into two sub-tasks, namely: (a) $\mathrm{U}^{c}\mathrm{CDR}$, where the query domain $q\in\{1,...,K\}$ but the classes are held out, i.e. $\mathcal{C_{\text{train}}}\cap\mathcal{C_{\text{test}}}\equiv\phi$. (b) $\mathrm{U}^{d}\mathrm{CDR}$, where the query domain $q\notin\{1,...,K\}$ but the classes are seen, i.e. $\mathcal{C_{\text{train}}}\equiv\mathcal{C_{\text{test}}}$. The goal of UCDR is to learn a mapping function that can place query and gallery images into a latent domain-independent subspace $\Phi$ , where images of same label but different domains are clustered together. The most challenging part is images from unseen domains and unknown classes also need to be mapped correctly in $\Phi$.
CLIP. Contrastive Language-Image Pre-training (CLIP) [31] is a multi-modality model pre-trained from 400 million image-text pairs. It aligns texts and images by an image encoder $f(\cdot)$ and a text encoder $g(\cdot)$. CLIP classifies images in a zero-shot way, based on the similarity between image features $f(\boldsymbol{x})$ and text features. The text features of various class-wise caption $\left\{\boldsymbol{t}_{i}\right\}_{i=1}^{|C_{train}|}$ can be formulated as $\left\{g(\boldsymbol{t}_{i})\right\}_{i=1}^{|C_{train}|}$, which are produced manually by filling class names into a text template. For example, given a text template “a photo of a [CLASS].”, then exchange [CLASS] with class names to construct class-wise caption, e.g., [CLASS] $\rightarrow$ cat / dog / bike. Formally speaking, given an image $\boldsymbol{x}$ and class caption $\left\{\boldsymbol{t}_{i}\right\}_{i=1}^{|C_{train}|}$ , CLIP output a prediction by:
[…]
where $\otimes$ is cosine similarity. Benefiting from large-scale pre-training and alignment of images and text, CLIP is robust for various visual appearance changes [15], which is suitable for UCDR task. However, fine-tuning CLIP will distort pre-trained features and lead to poor out-of-distribution performance [19]. Therefore, how to effectively leverage this powerful foundation model for UCDR becomes an important problem.
Prompt Tuning. Compared with full fine-tuning, prompt tuning prevents pre-trained feature distortion by introducing task-specific trainable parameters termed prompts into the input while keeping pre-trained model frozen. Specifically, [41] incorporate learnable text prompts $\mathbf{P_{t}}=\left\{p_{t}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{t}}$ into the text template as the input of CLIP text encoder, where $N_{t}$ and $\mathbb{R}^{\ell}$ indicate text prompts length and dimensions, respectively. Then the text template can be represented as “$\mathbf{P_{t}}$ a photo of a [CLASS]”. Furthermore, [17] proposed visual prompts for ViT which can be formulated as $\mathbf{P_{v}}=\left\{p_{v}^{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{N_{v}}$, where $N_{v}$ means the number of visual prompts and $p_{v}$ is a learnable parameter with $\ell$ dimension. Then the visual prompts are inserted into the first Transformer layer, named VPT-Shallow in [17] :
${\left[\mathbf{x}_{1},\mathbf{Z}_{1},\mathbf{E}_{1}\right]}$ $=L_{1}\left(\left[\mathbf{x}_{0},\mathbf{P_{v}},\mathbf{E}_{0}\right]\right),$ (2)
${\left[\mathbf{x}_{i},\mathbf{Z}_{i},\mathbf{E}_{i}\right]}$ $=L_{i}\left(\left[\mathbf{x}_{i-1},\mathbf{Z}_{i-1},\mathbf{E}_{i-1}\right]\right)\quad i=2,3,\ldots,N_{L},$
$\mathbf{Output}$ $=\operatorname{Head}\left(\mathbf{x}_{N_{L}}\right),$
where $\mathbf{Z}_{i}\in\mathbb{R}^{N_{v}\times\ell}$ represents the prompt embeddings computed by the $i^{th}$ Transformer layer in the CLIP image encoder, $\mathbf{E}_{0}$ means the image patch embeddings extracted from a CNN feature map [8]. In addition, $\mathbf{x_{0}}$ is CLS token, which produces $\mathbf{y}$, and served as the final output of image representation.
3.2 Prompt-to-Simulate
Prompt Units Learning (PUL). To capture domain-specific and class-specific knowledge in PUL, we introduce two types of prompt units: domain prompt units $DP$ and semantic prompt units $SP$, as shown in Fig. 2 (left). Specifically, we employ $K$ learnable vectors as domain prompt units to represent $K$ source domains in training datasets:
$DP=\left\{dp_{i}\in\mathbb{R}^{\ell}\right\}_{i=1}^{K}.$ (3)
Similarly, we introduce a set of learnable semantic prompts for $\left|\mathcal{C_{\text{train}}}\right|$ seen classes which can be formulated as:
$SP=\left\{sp_{i}\in\mathbb{R}^{\ell}\right\}^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}.$ (4)
To ensure that each prompt only learns knowledge related to a specific domain or class, we iteratively update single $dp$ and $sp$ w.r.t. an input image. To achieve this, we mask the irrelevant domain and semantic prompt units by $\mathbf{M_{id}}=\left\{\mathbf{m_{id}}^{i}\in[0,1]\right\}^{K}_{i=1}$ and $\mathbf{M_{is}}=\left\{\mathbf{m_{is}}^{i}\in[0,1]\right\}^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}$. Assume that, the input image is from domain $\mathbf{d}$ and class $\mathbf{c}$, than $\mathbf{m_{id}^{d}}$ and $\mathbf{m_{is}^{c}}$ are $1$ and other entries are 00. After mask operation, two masked prompt units are then concatenated with other input tokens to formulate a prompted input:
$\mathbf{x_{p}}=[\mathbf{x_{0}},\mathbf{M_{id}}\odot DP,\mathbf{M_{is}}\odot SP,\mathbf{E_{0}}],$ (5)
where $\odot$ discards prompts with corresponding mask of 00 preserves others of $1$. Correspondingly, image feature is obtained in the same way as Eq. 2. Since text feature can be considered as the class prototype, we introduce the learnable text prompts $\mathbf{P_{t}}$ as we discussed in Sec. 3.1 to achieve more flexible training. Additionally, we design a new text template: “a photo of $[CLASS]^{i}$ from $\mathbf{P_{t}}$ domain.”, to better suit the UCDR task. Given the prompted input $\mathbf{x_{p}}$ and the class caption $\left\{\boldsymbol{t}_{i}\right\}_{i=1}^{|C_{train}|}$ generated by above text template, the training loss can be formulated as:
$\mathcal{L}=-\sum^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}y_{i}\log{(f(\mathbf{\mathbf{x_{p}}})\otimes g\left(\boldsymbol{t}_{i}\right))},$ (6)
where $\otimes$ is cosine similarity and $y$ indicates the class label. The above procedure results in a mask-and-align objective. Our masked input ensures prompts to capture specific source knowledge from domains and categories. Then, the image features and text features are aligned among different domains in a union semantic space.
Context-aware Simulator Learning (CSL). In CSL, our goal is to train a module termed Content-aware Prompt Simulator (CaPS) to generate two Content-aware Dynamic Prompts (CaDP) similar to Prompt Units which can well impact CLIP to extract generalized features of unseen samples. Towards this goal, we concatenate two sets of Prompt Units with image patches as CaPS’s input, which can be formulate as $[PT_{d},PT_{s},DP,SP,\mathbf{E_{0}}]$, where $PT_{d}\in\mathbb{R}^{\ell},\;PT_{s}\in\mathbb{R}^{\ell}$ are domain and semantic prompt template for CaDP generation and $\mathbf{E_{0}}$ are unseen sample’s embedding.
Inspired by the ieda of meta-learning, we train CaPS under a simulated test scenario by a relevant mask operation. Specifically, we mask relevant domain and semantic prompt units w.r.t. input samples by $\mathbf{M_{rd}}=\left\{\mathbf{m_{rd}}^{i}\in[0,1]\right\}^{K}_{i=1}$ and $\mathbf{M_{rs}}=\left\{\mathbf{m_{rs}}^{i}\in[0,1]\right\}^{\left|\mathcal{C_{\text{train}}}\right|}_{i=1}$, where $\mathbf{m_{rd}^{d}}=0$, $\mathbf{m_{rs}^{c}}=0$ and others are $1$ if the input sample is from domain $\mathbf{d}$ and class $\mathbf{c}$. Then we can hide the prompt units corresponding to input samples by multiplying them with $\mathbf{M_{rs}}$ and $\mathbf{M_{rs}}$ mask matrices. The whole process can be formulated as:
${\left[\mathcal{P}_{d},\mathcal{P}_{s}\right]}$ $=\mathcal{M}([PT_{d},PT_{s},\mathbf{M_{rd}}\odot DP,\mathbf{M_{rs}}\odot SP,\mathbf{E_{0}}]),$ (7)
$\mathbf{x_{p}}$ $=[\mathbf{x_{0}},\mathcal{P}_{d},\mathcal{P}_{s},\mathbf{E_{0}}],$
Similarly, we use the same objective as Eq. 6 to train the newly introduced $PT_{d},PT_{s}$ and $\mathcal{M}$ while remaining part would be fixed in this stage.
3.3 Retrieval by ProS
After two stage training, now we can generate CaDP by CaPS, and extract image features by CLIP with the help of CaPS.Fig. 2 (right) shows how to employ our proposed ProS for retrieval. Since we only need image features for retrieval, CLIP text encoder $g$ is discarded. We first use CaPS $\mathcal{M}$ to generate two CaDP, and then use CLIP Image Encoder $f$ to get well generalized image features $\mathbf{y}$:
${\left[\mathcal{P}_{d},\mathcal{P}_{s}\right]}$ $=\mathcal{M}([PT_{d},PT_{s},DP,SP,\mathbf{E_{0}}]),$ (8)
$\mathbf{x}$ $=f\left(\left[\mathbf{x_{0}},\mathcal{P}_{d},\mathcal{P}_{s},\mathbf{E}_{0}\right]\right),$
$\mathbf{y}$ $=\operatorname{Head}\left(\mathbf{x}\right).$
To perform retrieval, all images in $\mathcal{C}_{\text{test}}$ are indexed by Eq. 8 to formulate the gallery. Then, given a query, we do the same thing to obtain the query feature and produce a rank list by sorting similarities between query and all features in gallery. In the next section, we would give a comprehensive study about retrieval results of our proposed ProS.
Query Domain Method UCDR $\mathrm{U^{d}}\mathrm{CDR}$
Unseen Gallery Mixed Gallery mAP@200 Prec@200
mAP@200 Prec@200 mAP@200 Prec@200
Sketch SnMpNet [29] 0.3007 0.2432 0.2624 0.2134 0.3529 0.1657
SASA [37] 0.5262 0.4468 0.4732 0.4025 0.5733 0.5290
CLIP-Full 0.5367 0.4666 0.4788 0.4136 0.6128 0.3806
CoOp [41] 0.5512 0.4947 0.4995 0.4479 0.6374 0.4245
VPT [17] 0.6216 0.5676 0.5609 0.5130 0.6769 0.4405
ProS (Ours) 0.6457 0.6001 0.5843 0.5463 0.7385 0.4911
Quickdraw SnMpNet [29] 0.1736 0.1284 0.1512 0.1111 0.1077 0.0509
SASA [37] 0.2564 0.1970 0.2116 0.1651 0.1805 0.1549
CLIP-Full 0.2011 0.1522 0.1622 0.1196 0.1820 0.0723
CoOp [41] 0.1484 0.1237 0.1183 0.0961 0.1834 0.084
VPT [17] 0.2467 0.2092 0.1953 0.1688 0.2367 0.0982
ProS (Ours) 0.2842 0.2544 0.2318 0.2127 0.2889 0.1186
Painting SnMpNet [29] 0.4031 0.3332 0.3635 0.3019 0.4808 0.4424
SASA [37] 0.5898 0.5188 0.5463 0.4804 0.5596 0.5178
CLIP-Full 0.6558 0.5926 0.6083 0.5478 0.6189 0.3688
CoOp [41] 0.6886 0.6207 0.6509 0.5884 0.6625 0.4128
VPT [17] 0.7138 0.6503 0.6752 0.6153 0.6618 0.4105
ProS (Ours) 0.7516 0.6955 0.7120 0.6612 0.7227 0.4615
Infograph SnMpNet [29] 0.2079 0.1717 0.1800 0.1496 0.1957 0.1764
SASA [37] 0.2823 0.2425 0.2491 0.2113 0.2340 0.2093
CLIP-Full 0.5332 0.4893 0.4718 0.4309 0.5311 0.3330
CoOp [41] 0.5285 0.4807 0.4820 0.4390 0.5530 0.3546
VPT [17] 0.5434 0.4957 0.4870 0.4468 0.5690 0.3566
ProS (Ours) 0.5798 0.5442 0.5219 0.4956 0.6056 0.3962
Clipart SnMpNet [29] 0.4198 0.3323 0.3765 0.2959 0.552 0.5074
SASA [37] 0.4397 0.3670 0.3940 0.3295 0.684 0.6361
CLIP-Full 0.6880 0.6200 0.6423 0.5755 0.6922 0.4174
CoOp [41] 0.7025 0.6414 0.6648 0.6068 0.7495 0.4776
VPT [17] 0.7344 0.6785 0.6942 0.6409 0.7536 0.4770
ProS (Ours) 0.7648 0.7186 0.7228 0.6815 0.8105 0.5298
Average SnMpNet [29] 0.3010 0.2418 0.2667 0.2144 0.3378 0.2686
SASA [37] 0.4189 0.3544 0.3748 0.3178 0.4463 0.4094
CLIP-Full 0.5229 0.4641 0.4727 0.4175 0.5274 0.3144
CoOp [41] 0.5238 0.4722 0.4831 0.4356 0.5572 0.3507
VPT [17] 0.5720 0.5203 0.5225 0.4770 0.5796 0.3566
ProS (Ours) 0.6052 0.5626 0.5546 0.5195 0.6332 0.3994
Table 1: UCDR and $\mathrm{U^{d}CDR}$ evaluation results on DomainNet. UCDR has two different gallery settings, i.e. the gallery set consists of (1) only unseen class images (Unseen Gallery) or (2) both seen and unseen images from Real domain (Mixed Gallery).
4 Experiment
4.1 Experimental setting
Datasets. Following the existing works [29, 37], we evaluate the effectiveness of our ProS on three popular datasets, including DomainNet [30], Sketchy [35, 26], and TU-Berlin [9, 26]) under three different cross-domain retrieval settings, i.e., UCDR, $\mathrm{U}^{c}\mathrm{CDR}$ and $\mathrm{U}^{d}\mathrm{CDR}$ as we introduced in Sec. 3.1.
DomainNet [30] contains 596,006 images collected in 6 domains (Real, Sketch, Quickdraw, Infograph, Clipart, Painting) from 345 categories and split into 245, 55, 45 categories for training, validation and testing respectively. To satisfy the unseen domain requirement in both $\mathrm{U^{d}CDR}$ and $\mathrm{UCDR}$ protocol, leave-one-out evaluation protocol is applied, where we iteratively select one domain as unseen query set and use other domains for training. Additionally, for $\mathrm{U^{d}CDR}$ evaluation, we select 45 categories from training set and choose $25\%$ samples from unseen domains as queries (except for Quickdraw, which is $10\%$ due to it being too large). The gallery set is constructed with Real images from unseen categories or mixed with seen categories, termed Unseen Gallery and Mixed Gallery respectively.
Sketchy [35, 26] contains 75,471 Sketches and 73,002 images from 125 categories where the train, validation, and test splits contain 93 , 11 and 21 categories [37, 29], respectively. TU-Berlin [9, 26] has 20,000 Sketches and 204,489 images which are separated into 200 categories for training, 20 categories for validation, and 30 categories for testing [37, 29]. Both Sketchy and TU-Berlin are used for $\mathrm{U^{c}CDR}$.
Evaluation Metrics. For fair comparison, we utilize the same evaluation metrics following [29, 37]. For Sketchy and DomainNet, we evaluate precision and mean Average Precision w.r.t. top-200 candidates (Prec@200 and mAP@200). For TU-Berlin, we use Prec@100 and mAP@all instead as evaluation metrics.
Implementation Details. In all our experiments, we use a fixed image encoder and text encoder initialized with pre-trained CLIP (ViT-B/32). For Context-aware Prompt Simulator $\mathcal{M}$, we adopt a two layer ViT with random initialization. Following  [41], the length of text prompt $\mathbf{P_{t}}$ is set to 16 with 512 dimension. In addition, the dimension of other prompts including prompt units and Content-aware Dynamic Prompts are set to 768. The training epochs are set to 10 with early stopping of 2 epochs based on the evaluation performance for all the datasets with a batch size of 50. We use Adam optimizer with a learning rate of 1e-3 and adopt a cosine decay strategy for the learning rate.
4.2 Main Results
Baselines. We compare our ProS with two groups of methods, including two existing UCDR methods, i.e., SnMpNet [29] which adopt a ResNet as backbone and SASA [37], which implement with a ViT. For fair evaluation, we further build three CLIP-based baselines, including CLIP-full and two prompt tuning methods, i.e., CoOp [41] and VPT [17] where CLIP-full indicate fine-tunes full CLIP.
Results on UCDR and $\mathbf{U^{d}CDR}$. We firstly evaluate our ProS methods on DomainNet under UCDR and $\mathrm{U^{d}}\mathrm{CDR}$ settings. The UCDR results in Tab. 1 shows that our ProS method outperforms all baselines across all domains, demonstrating the superiority of ProS in terms of generalization on unseen domains and categories. Furthermore, we have highlight the following conclusions: First (CLIP vs. ViT): all CLIP-based methods outperform ViT-based method which prove our assumption that general knowledge gained in large-scale pre-training can enhance the model’s ability to handle the UCDR task. Specifically, CLIP-Full achieves 9.79% mAP@200 improvement than SASA. Second (Prompt Tuning vs. Fine-Tuning): VPT and CoOp all both outperforms CLIP-Full, demonstrating the effectiveness of prompt tuning to adapt CLIP. Third (ProS vs. Prompt Tuning): compared with VPT, our ProS further improves the performance by 3.21% mAP@200 on average, indicating that ProS effectively apply prompt tuning for UCDR task.
From reported $\mathbf{U^{d}CDR}$ results in Tab. 1, we can draw the conclusion that our ProS consistently outperforms all competitors in terms of cross-domain alignment and generalization. Specifically, comparing with average results, our method surpasses VPT by $5.36\%$ in mAP@200.
Method Sketchy TU-Berlin
mAP@200 Prec@200 mAP@all Prec@100
SnMpNet [29] 0.5781 0.5155 0.3568 0.5226
SASA [37] 0.6910 0.6090 0.4715 0.6682
CLIP-Full 0.6553 0.6145 0.6076 0.7158
CoOp [41] 0.5074 0.4659 0.5585 0.6759
VPT [17] 0.6588 0.6105 0.5574 0.6815
ProS (Ours) 0.6991 0.6545 0.6675 0.7442
Table 2: $\mathrm{U^{c}CDR}$ evaluation results on Sketchy and TU-Berlin. Consistent with [29, 37], we use Prec@200 and mAP@200 in Sketchy, and Prec@100 and mAP@all in TU-Berlin as evaluation metrics.
Results on $\mathbf{U^{c}CDR}$. In Tab. 2, we explore the ability of ProS in handling semantic shift under $\mathrm{U^{c}CDR}$ setting on Sketchy and TU-Berlin. We observe that: First, our ProS shows consistent improvement compared to CLIP-Full and VPT, revealing that our ProS can enhance the ability of CLIP to manage semantic shifts and generalize to unseen categories. Second, compared to SASA, our ProS has a larger improvement on TU-Berlin than that on Sketchy($19.6\%$ vs. $0.81\%$ in mAP@200). We could see CLIP, CoOp and VPT, all of which use CLIP as the backbone, exhibit relatively low mAP on Sketchy, potentially caused by a poor performance of the visual backbone when handling this dataset. However, our method still surpasses all competitor and achieves the best performance.
4.3 Ablation Study
In this section, we conduct extensive ablation experiments to reveal the contributions of each component. For all ablations, we select Infograph as unseen query set from DomainNet and evaluate ProS under UCDR setting with two gallery configurations.
Impact of Each Components. To validate the effectiveness of each component in our ProS, we individually remove each component and experiment with the following five settings: 1) “w/o $SP$” (remove Semantic Prompt Units). 2) “w/o” $DP$ (remove Domain Prompt Units). 3) “w/o Mask” (remove mask operation in training). 4) “w/o $\mathcal{M}$” (remove Content-aware Prompt Simulator). 5) “w/o CLS” (remove CLS token). The experiment results are shown in Tab. 3. We observe that: First, removing either $SP$ or $DP$ led to a performance degradation of -2.65% and -2.56% in mAP@200 under mixed gallery settings, indicating their essential roles in Content-aware Prompt Simulation stage. Second, without the mask operation, mAP@200 has decreased by 1.27% which proves the effectiveness of mask operation. Third, when we remove CaPS $\mathcal{M}$ from our model, it lead to worse performance than full ProS, highlighting the significant benefits of $\mathcal{M}$ for UCDR task. Forth, training $CLS$ token is absolutely necessary as its omission caused a performance reduction of 6.18
Method Unseen Gallery Mixed Gallery
mAP@200 Prec200 mAP200 Prec200
ProS 0.5798 0.5442 0.5219 0.4956
w/o $SP$ 0.5529 0.5142 0.4954 0.4641
w/o $DP$ 0.5566 0.5190 0.4963 0.4680
w/o Mask 0.5692 0.5324 0.5092 0.4809
w/o $\mathcal{M}$ 0.5573 0.5128 0.5053 0.4685
w/o $CLS$ 0.5241 0.4846 0.4601 0.4260
Table 3: Different component evaluated on DomainNet under UCDR protocol with Infograph as the unseen domain for queries.
Impact of Transformer Layer in $\mathcal{M}$. We further investigate the impact of Transformer layers in Context-aware Prompt Simulator $\mathcal{M}$. From the Tab. 4, it can be found that: 2 layer ViT yield the best performance. Although employing only 1 layer can slightly reduce computational costs, it results in a worse performance. Hence, for an optimal trade-off between accuracy and efficiency, a 2 layer ViT is chosen to implement $\mathcal{M}$.
$\mathcal{M}$’s Layer Unseen Gallery Mixed Gallery
mAP@200 Prec@200 mAP@200 Prec@200
1 0.5752 0.5379 0.5177 0.4900
2 0.5798 0.5442 0.5219 0.4956
3 0.5790 0.5429 0.5210 0.4942
6 0.5743 0.5388 0.5160 0.4893
12 0.5598 0.5228 0.5036 0.4742
Table 4: Ablation study on effectiveness of different transformer layer in Context-aware Prompt Simulator $\mathcal{M}$.
Figure 3: Evaluation results of two prompts length. (a) investigate the impact of text prompt length. (b) analyse Content-aware Dynamic Prompt length generated by CaPS $\mathcal{M}$,where 0-0 represent VPT and 1-1 means one CaDP for domain and one for semantic.
[…]
Figure 5: Retrieval Results under UCDR protocols on DomainNet. (a) displays the retrieval results of “Peas” by the query from a unseen query domain. (b) shows the retrieval results of a few queries from Quickdraw. True positives and false positives are shown with green and red borders, respectively.
Impact of two prompts length. We further conduct experiments to analyse the impact of two key prompt length in ProS. i.e.,text prompts and Content-aware Dyanmic Prompt (CaDP). Fig. 3(a) shows that when text prompts length increases, the performance improves to its peak when length is 16 and then gradually decreases. Therefore, we set the text prompt length to 16 for all datasets. In Fig. 3(b), we find that ProS achieve the state-of-the-art with only 1-1 CaDP. We speculate that this might be consistent with Prompt Units, which also use a single prompt to represent one domain or category.
4.4 Qualitative Analysis
[…]
To further demonstrate the effectiveness of our ProS, we visualize a few retrieval results produced by ProS. Fig. 5 (a) shows top five retrieved candidates of peas with query images from different domains, while Fig. 5 (b) shows top ten retrieved candidates with queries from Quickdraw. From the figure we can see that Sketch and Quickdraw images lack enough information such as color, texture, and details while Inforgraph has a object co-occurrence problem. They all obtain worse performance compared with other domains. This phenomenon will be exacerbated when the gallery contains both seen and unseen classes due to the interference caused by seen classes. Towards this phenomenon, we can conclude that: due to the quality issues of the dataset, the performance of our APL in certain domains may be relatively poor.
5 Conclusion
In this paper, to the best of our knowledge, we take the first attempt to deploy CLIP for Universal Cross-Domain Retrieval (UCDR) with a prompt tuning paradigm and propose a more generalized prompt method named Prompt-to-Simulate (ProS). Specifically,ProS can dynamically fits unknown domain and category distribution with guidance of source knowledge by a two-stage training paradigms following a mask-and-align objective. With the above approach, our model obtains quite strong performance under UCDR compared to state-of-the-art.